OpenStudy (bananas1234):

Find k(3) for the function k(n) = 4n^2

OpenStudy (mhchen):

Just plug 3 in so k(3) = 4(3^2) Can you evaluate that?

OpenStudy (bananas1234):

24?

OpenStudy (mhchen):

3^2 = 3 * 3 = 9 so 4(9) = ?

OpenStudy (bananas1234):

i mean 36

OpenStudy (mhchen):

ye

OpenStudy (bananas1234):

thanks

OpenStudy (mhchen):

np